Nov 13, 1999 OWL shift summary

Richard Kessler (rkessler@fnal.gov)
Sat, 13 Nov 1999 07:05:03 -0600 (CST)

Shift Summary for OWL shift, Nov 13, 1999

On Shift: R.Kessler, M.Timko

Hours of beam available at all intensities: 5.8
Hours of beam available at nominal intensity: 5.8
Hours of E832 data (special runs not included): 5.7
Hours of special runs (laser scans, muon runs...):

========================================================================
Accounting of lost time: tape change.
========================================================================
Runs ended this shift: Run 14956 - 92K 3pi0 D
========================================================================
Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: None
========================================================================
Overview of shift: E799 physics data.
========================================================================
Description of Shift:

Smooth data taking with a few minor glites:

- there was a pre-mature run abortion, "Boy I'm full",
even though the DAQ monitor showed the tapes to be
just 70% full. Experts contacted to investigate.

- The laser illum plot was EMPTY and no Val-audio alarms.
It was traced to the laser trigger being turned off.
I turned it back on at the ktrig window and that worked fine.
So there is no laser data for run 14957 for a few spills
near spill 80.

- FPED got stuck on-spill for two spills in a row.

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift: none
========================================================================
Items to Follow up on Next Shift: none